OpenStudy (anonymous):

multiplying radical expression. simplify. 3 squared 5 x 18 squared ???

OpenStudy (anonymous):

\[3^2*5(18)^2\] is this your problem?

OpenStudy (anonymous):

no the squared is a radical sign

OpenStudy (anonymous):

\[3\sqrt{5}*\sqrt{18}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok well, this first thing we do is realize that 18 can be written as 9 times 2 so we get this: \[3\sqrt{5}*\sqrt{9*}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[3\sqrt{5}*\sqrt{9*2}\]

OpenStudy (anonymous):

The 9 comes out of the radical as 3, the two stays underneath the radical, so we get: \[3\sqrt{5}*3\sqrt{2}\]

OpenStudy (anonymous):

That is as far as we can go.

OpenStudy (anonymous):

okay thank you :)